How would I turn the point slope form linear equation, y-3=3(X+1), into standard form? Also, how would I turn the slope intercept form equation, y=3x+-3, into standard form?
Answer by rothauserc(4718)

The standard form for a line is
:
Ax +By = C, where A is a positive integer and B and C are integers
:
1) y-3 = 3 (x+1)
y-3 = 3x+3
3x -y = -6
:
2) y = 3x+-3
3x -y = 3
: